Please can you help me find the intersection points (x,y)for each of the following problems...
1)y=-2x+9
2)y=-1x+8
3)y=3/4x+1/2
4)y=-2x+3

Answer by bucky(2189)

To find the intersection points, follow these two rules:
.
To find the x-axis intersection point, set y equal to zero and solve for x.
.
To find the y-axis intersection point, set x equal to zero and solve for y.
.
Let's now apply those rules:
.
Given: y=-2x+9
.
Set y equal to zero and this equation becomes 0 = -2x + 9. Subtract 9 from both sides and
the equation becomes -9 = -2x. Solve for x by dividing both sides by -2 and you get the
answer: x = -9/-2 = 9/2. The graph crosses the x-axis at the point where x = 9/2.
.
Set x equal to zero and solve for y. When you set x equal to zero the equation becomes
y = 0 + 9. That's the answer. The graph crosses the y-axis at the point where y = +9.
.
Next problem. Given: y = -1x + 8
.
Set y equal to zero and the equation becomes 0 = -1x + 8. You can solve for x by adding
1x (or just x) to both sides of the equation. When you do, the equation becomes
.
x = +8
.
That's the answer. The graph crosses the x axis at the point where x = +8.
.
Next, set x equal to zero and the equation becomes y = 0 + 8 or just y = +8. And that's the
other answer. The graph crosses the y-axis at the point where y = +8.
.
Third problem. Given: y = (3/4)x + 1/2
.
Set y = 0 and the equation then is 0 = (3/4)x + 1/2. You can get rid of the denominators
by multiplying both sides of the equation (all terms) by 4. When you do the equation
then becomes: 0 = 3x + 2. Subtract 2 from both sides and the equation then becomes
-2 = 3x. Solve for x by dividing both sides of the equation by 3 to arrive at the solution
of x = -2/3. So the graph crosses the x-axis at the point where x = -2/3.
.
Then set x equal to zero and the equation reduces to y = 0 + 1/2. This is the answer.
The graph crosses the y-axis at the point where y = +1/2.
.
Last problem:
.
Given: y = -2x + 3
.
Find the point where the graph crosses the x-axis by first setting y equal to zero. When
you do that, the equation becomes 0 = -2x + 3. Subtract 3 from both sides and the equation
changes to -3 = -2x. Solve for x by dividing both sides by -2 and you then find that
x = -3/-2 = 3/2. The graph crosses the x-axis at the point where x = 3/2.
.
Then find the point where the graph crosses the y-axis by setting x equal to zero. When
you do that the equation becomes y = 0 + 3. That's the answer. The graph crosses the y-axis
at the point where the value of y is +3.
.
Hope this helps you to understand how to do these problems.